1) Solve the quadratic equations by factoring: 5x² = 125
2) Solve the quadratic equation using the square root property: x² - 49 = 0
3) Solve the quadratic equation by completing the square: x² - 8x = 7
4) Solve the quadratic equation using the quadratic formula: 8x² - 10x - 2 = 0
5) Solve the quadratic equation using the quadratic formula: 2x² - 5 = 4
6. Solve using the quadratic formula:
x² – 3x – 3 = -5
please if you can helpm me. thank you

Answer by checkley71(8403)

I'LL SOLVE THE FIRST 3 FOR YOU
5X^2=125
5X^2-125=0
5(X-5)(X+5)=0
X-5=0
X=5 SOLUTION
X+5=0
X=-5 SOLUTION
------------------------------
X^2-49=0
(X-7)(X+70=0
X-7=0
X=7 SOLUTION
X+7=0
X=-7 SOLUTION
---------------------------
X^2-8X=7
X^2+8X+16=7+16
(X+4)(X+4)=23
X+4=SQRT23
X+4=4.7958
X=4.7958-4
X=.7958